A popular exhibition among riders in the Philippine is the “banking sa curvada” which is a feat that takes advantage of the structure of the curved road. Originally built for race tracks, the banking is adopted in curved national roads and highways to maximize tire-road friction which keeps the vehicle from tumbling out of the road. Suppose a curvada with a radius of 150m is banked at the right angle for a maximum speed of 25 m/s. At this speed, what is the coefficient of friction between the road and the tire?


1
Expert's answer
2021-05-31T15:14:34-0400

Gives

V=25m/sec


Radius

R=150 m

Friction Cofficient ="\mu"



We know that

Curved road obtained maximum speed

"{V_{max}}^2=\mu Rg"

"\mu=\frac {V^2_{max}}{Rg}"

Put value

"\mu=\frac{{25}^2}{150\times 9.8}"

"\mu=0.425"